OpenStudy (anonymous):

You place a cup of 200ºF coffee on a table in a room that is 67ºF, and 10 minutes later, it is 195ºF. Approximately how long will it be before the coffee is 180ºF? Use Newton's law of cooling: T(t)=T(subscriptA)+(T[subscripto]−T[subscriptA])e^(−kt) A. 40 minutes B. 15 minutes C. 1 hour D. 35 minutes

OpenStudy (anonymous):

ah back to newton's law yes?

OpenStudy (anonymous):

wanna do it the easy way?

OpenStudy (anonymous):

T(t) = final temperature, T(a) = ambient temp (that is, the temperature of the room), T(to) = original temperature, k = constant you need to solve for, t = time

OpenStudy (anonymous):

just plug the the variables from the first sentence into the equation, solve for k, and then plug the variables with the k value you calculated, only this time use 180 instead of 195 as the final temperature, and solve for t this time

OpenStudy (anonymous):

you understand all that or you want to do it step by step or you want the answer?

OpenStudy (anonymous):

the answer would be nice lol

OpenStudy (anonymous):

give me as second i am going to do it the easy way, then give you the answer

OpenStudy (anonymous):

i get 42 so i guess 40 is closest

OpenStudy (anonymous):

let me check calculator again

OpenStudy (anonymous):

yup. that is what i get. you want easy method?

OpenStudy (anonymous):

thanks, and sure

OpenStudy (anonymous):

ok first off we need to work with the differences in the temp, not the actual temp. don't let this throw you off 200- 67 = 133 195-67 = 128 185 - 67 = 113

OpenStudy (anonymous):

those are the numbers we will work with on the right.

OpenStudy (anonymous):

decreased from 133 to 128 in ten minutes formula is \[133(\frac{128}{133})^{\frac{t}{10}}\]

OpenStudy (anonymous):

you see how i put the numbers in. the ten is from the ten minutes, the initial values is 133 and then end value is 128

OpenStudy (anonymous):

then solve \[113=133(\frac{128}{133})^{\frac{t}{10}}\] for t and if that is hard just do it here http://www.wolframalpha.com/input/?i=113%3D133%28128%2F133%29^%28x%2F10%29

OpenStudy (anonymous):

cool, thanks